# Comprehensive X-ray Observations of the Exceptional Ultra-long X-ray and Gamma-ray Transient GRB 250702B with Swift, NuSTAR, and Chandra: Insights from the X-ray Afterglow Properties

### v1

@@ 6 Conclusions
We demonstrate that the overall characteristics of GRB 250702B do not cleanly match either known relativistic TDEs or ultra-long GRBs. With the exception of the initial soft X-ray detection a full day before the initial gamma-ray triggers, the properties have a closer match to those seen by ultra-long GRBs, but require an even longer lived central engine ($\gtrsim$33 days). In particular, the soft X-ray precursor (H. Q. Cheng et al., 2025a) is not easily explained in the standard collapsar model (Woosley1993). As such, we favor a hybrid stellar-mass black hole progenitor, such as a micro-TDE (P. Beniamini et al., 2025), though see also Neights2025 and J. Granot (2025) for alternative explanations.

### v2

@@ 6 Conclusions
We demonstrate that the overall characteristics of GRB 250702B do not cleanly match either known relativistic TDEs or ultra-long GRBs. With the exception of the initial soft X-ray detection a full day before the initial gamma-ray triggers, the properties have a closer match to those seen by ultra-long GRBs, but require an even longer lived central engine ($\gtrsim$33 days). In particular, the early soft X-ray detections (H. Q. Cheng et al., 2025a; EP250702a-arxiv) are not easily explained in the standard collapsar model (Woosley1993). As such, we favor a hybrid stellar-mass black hole progenitor, such as a micro-TDE (P. Beniamini et al., 2025), though see also Neights2025 and J. Granot et al. (2025) for alternative explanations.